### v1

@@ 1 Introduction
The difference is subtle but crucial. When simulated, agents operate in a virtual sandbox environment designed to imitate the dynamics of the target real-world deployment in which perceptions (e.g., data from sensors, the flow of time) and actions (e.g., agents’ movements in the environment) are controlled by the simulator, and can be reproduced deterministically. When deployed, agents are attached to specific hardware and runtimes, communicate through networks, and are tied to the flow of real time. Agents could either run in virtual environments as concurrent (and often distributed) processes interacting with shared resources (e.g., files, databases, network sockets) hence perceiving the state of such resources in real time and acting on them. In some cases, agents may even be embodied in physical entities (e.g., robots, drones, embedded devices) hence perceiving and acting through physical sensors and actuators. Any of these elements may introduce non-determinism and unpredictability in the system, to which the MAS must be ready to react.
@@
Finally, we exercise the prototype in a (simulated) multi-drone coordination scenario, to exemplify how it can be exploited as means for the early testing of BDI systems. More specifically, we show how simulating the same BDI specification with different performance/fidelity trade-offs may help in spotting issues that would have been otherwise overlooked.
@@ 2 Background
Here, we recall the main notions behind BDI agents and computer simulation, under a unifying perspective rooted in the notion of event. Accordingly, in Section˜2.1 we first summarise the state of the art of BDI agents architectures, then we analyse which events are involved in their lifecycle. We also report in section Section˜2.2 major definitions such as discrete-event, -time, and multi-agent-based simulation, to ground our selection for the mapping of BDI in simulation.
@@ 2.1 The BDI Model
The execution of a BDI agent boils essentially down to scheduling control-loop phases and handle events. Practically, this may vary across implementations and be fairly different between simulations and real-world executions. Specifically, defining how the control loop is scheduled may lead to different emerging behaviours of the overall MAS due to which interleaving of agent’s control loops is allowed by the implementation. On the one side, in simulations, the simulator is in charge of (reproducibly) virtualizing the flow of time. In this case, ‘executing the BDI system’ means to advance the simulation time while deciding which agent(s) shall execute their control-loop phases next. The goal here is emulate concurrency, while keeping the execution of the agents deterministic and reproducible. On the other side, in real-world deployments, the agents are executed by the operating system, the flow of time corresponds to the real-world time, and the agents are scheduled by the operating system’s scheduler, which may introduce non-determinism in the execution of the agents. In this case, the goal is to balance agents execution onto the available computational resources, and to exploit parallelism accordingly.
@@ 3 BDI Agents Simulation Methods
This seamless switching is the primary goal for our work [DBLP:conf/dsrt/Baiardi24]. In fact, keeping the same agent specification ensures that simulation results are not artefacts induced by the software translation into a simulation-compatible model. Additionally, if the specification is the same, issues discovered at deployment time that were not evident during testing will be at most imputable to the abstraction gap of the environment model. A recent work has tackled this problem for non-BDI agents, [Briola2025], by developing a simulation environment for the JADE MAS platform [DBLP:books/wi/BellifemineCG07]. The work focuses on achieving portability of the same agent behaviour between simulation and real-world execution, by implementing a simulation environment that mimics the JADE run-time environment. We exclude this work from our comparison, as JADE agents are not BDI-based.
@@ 4.1 Portable Agent Specifications
Platform-specific implementations – at least, one for real-world deployments and another for simulation – must be provided for such a platform-agnostic API. We refer to these implementations as the interpreters of the BDI agent specification. Which are responsible for executing the agent code by relying on the underlying platform functionalities. The two interpreters may of course differ in their implementations, yet both should guarantee the AgentSpeak(L) semantics is preserved. In this regard, a good practice is to let the two interpreters share as much code as possible: ideally implementing the agent’s control loop in a shared library, which allows for plugging in different platform-specific functionalities.
@@
Agent behaviour may depend on the passage of time. Hence, agents may need to observe the current time, usually through the system’s clock in a real-world deployment. In simulation, however, the flow of time is virtualized and controlled by the simulator, and it usually does not correspond to real-world time. Thus, access to time-dependent functionalities (e.g., timestamps, timeouts, delays) must not be implemented by accessing system-level time primitives directly. Accessing the system’s clock, or using sleep functions must be avoided in the agent code, and instead the environment API must provide the necessary functionalities to access the current time or to schedule timeouts and delays. It will be then the responsibility of the integration with the simulation platform to provide a proper implementation of such functionalities upon the simulator’s notion of time, guaranteeing semantic correspondence with the real-world counterpart.
@@
Agents in a MAS often need to communicate with each other. To support inter-agent communication, the environment API must provide functionalities to send and receive messages. While real-world (distributed) deployments would rely on networking protocols (e.g., HTTP, MQTT, XMPP, etc.), simulators should take care of simulating the dispatch of messages to their recipient, hence emulating the role of telecommunication networks. Typically, in MAS the environment acts as a relay for messages, and messages might be subject to some sort of propagation mechanism. The way messages are propagated may depend on the scenario, for instance, taking into account the environment topology by limiting the delivery of messages to reachable destinations, or introducing some form of delay or loss in the delivery of messages to better match with the expected real-world environment dynamics and stress the properties of the system under different conditions. Again, failing to abstract communication functionalities would make impossible to simulate the system, as real-world networking protocols cannot be used in a simulated environment.
@@ 4.4 Tool Requirements for the Integration
Following the key features outlined in above sections, the tools to be adopted for the integration must provide:
@@
  • simulation model featuring locality, position, and communication channels, to support simulating agents situates in physical or logical spaces;
@@ 5.2 BDI MAS in Alchemist
Each agent’s lifecycle phases (i.e., sense, deliberate, and act) are mapped onto separate DES events (Alchemist’s Reactions). As introduced in Section˜5.1, each such event has a duration modelled through a TimeDistribution. To guarantee that each agent’s control loop phases are executed in order, we implemented a custom TimeDistribution, (JaktaTimeDistribution), composed of three sub-TimeDistributions, one for each phase of the control loop. JaktaTimeDistribution maintains an internal state which tracks the current phase of the agent, and executes the next phase according to its configured TimeDistribution. This design allows us to flexibly support different granularities:
@@
Agents can reason on the current device status, as information stored inside Nodes’ is reified as perceived knowledge from JaKtA agents. Information inside nodes is stored as key-value pairs (Molecules and associated Concentrations (c.f. Section˜5.1), that agents can manipulate through actions. Changes are then treated by the simulator as “regular” environment modificationa consequent to simulated events.
@@ 6.3 Experimental Setup
We ensured code portability by ensuring that all interactions with the environment are done through the JaKtA environment API, which we implemented both for the real-world deployment and for the Alchemist simulator.
6.3.1 Running on the Alchemist Simulator
We use the simulator to compare the execution of the same MAS logic with different granularities, namely, we consider AMA, ACLI, and ACLP (cf. Section˜4.3). It turns out that, when simulating coarse-grained granularity levels (namely, AMA and ACLI) the agents are able to maintain the formation, while the fine-grained ACLP granularity level reveals that this result is an artefact of the over-simplification of the BDI programs. On a real UAV-programming workflow, this would have revealed a flaw in the agent program, or highlighted the need for higher responsiveness (reasoning cycle frequency) to react to the dynamics of the environment.
@@ 6.3.1 Running on the Alchemist Simulator
6.3.2 Deployment on a Concurrent System
@@ 6.4 Results
6.4.1 Analysis of the Simulation Results
@@
6.4.2 Concurrent System Deployment Results
@@ 8.1 Achievement of the work objectives
A BDI framework must support modular execution models, abstracting platform-specific functionalities and encapsulating them into a deployment-agnostic API, that allows external scheduling of agents’ reasoning cycles. The framework should maintain a clean separation between agent logic and execution infrastructure, allowing the same specification to run both in simulation and on real-world hardware without modifications. Implementations should interpret the same agent logic differently. Core to the simulated deployment is the ability to execute BDI agents deterministically, to ensure reproducibility. The same is not true in real-world deployments where concurrency and asynchrony are inherent, and sometimes even desirable and exploitable. Therefore, the deployment-agnostic API should virtualise functionalities covering all potential sources of non-determinism, such as random number generation and external events.
@@
Our integration of JaKtA and Alchemist demonstrates that a BDI framework can be embedded within a DES engine by treating agent execution as a series of simulation-driven events. By leveraging modularity in both technologies, we enable BDI agents to operate in a simulated environment without requiring code modifications. The environment is modeled within the simulator, preserving interactions with agents while ensuring portability. The proposed approach ensures that the same BDI specification can be tested in simulation and seamlessly deployed to real-world hardware, reducing development overhead and minimizing the reality gap.
@@ Acknowledgements
This work has been partially supported by: “WOOD4.0 - Woodworking Machines for Industry 4.0”, (CUP E69J22007520009) Emilia-Romagna regional project, call 2022, art. 6 L.R. N. 14/2014; and by the projects funded by the European Commission under the NextGenerationEU programme: PNRR – M4C2 – Investment 1.3, Partenariato Esteso PE00000013 – “FAIR—Future Artificial Intelligence Research” – Spoke 8 “Pervasive AI” and PNRR M4C2, Investment 3.3 (DM 352/2022) in collaboration with AUSL Romagna - CUP J33C22001400009 "Digital Twins Ecosystems for the clinical, strategic and process governance in Healthcare".

### v2

@@ 1 Introduction
The difference is subtle but crucial. When simulated, agents operate in a virtual sandbox environment designed to imitate the dynamics of the target real-world deployment in which perceptions (e.g., data from sensors, the flow of time) and actions (e.g., agents’ movements in the environment) are controlled by the simulator, and can be reproduced deterministically. When deployed, agents are attached to specific hardware and software runtimes, communicate through networks, and are tied to the flow of real time. Agents could either run in virtual environments as concurrent (and often distributed) processes interacting with shared resources (e.g., files, databases, network sockets) hence perceiving the state of such resources in real time and acting on them. In some cases, agents may even be embodied in physical entities (e.g., robots, drones, embedded devices) hence perceiving and acting through physical sensors and actuators. Any of these elements may introduce non-determinism and unpredictability in the system, to which the MAS must be ready to react.
@@
Finally, we exercise the prototype in a (simulated) multi-drone coordination scenario, to exemplify how it can be exploited as means for the early testing of BDI systems. More specifically, we show how simulating the same BDI specification with different complexity/fidelity trade-offs may help in spotting issues that would have been otherwise overlooked.
@@ 2 Background
Here, we recall the main notions behind BDI agents and computer simulation, under a unifying perspective rooted in the notion of event. Accordingly, in Section˜2.1 we first summarise the state of the art of BDI agents architectures, then we analyse which events are involved in their lifecycle. We also report in Section˜2.2 major definitions such as discrete-event, -time, and multi-agent-based simulation, to ground our selection for the mapping of BDI in simulation.
@@ 2.1 The BDI Model
The execution of a BDI agent boils essentially down to scheduling control-loop phases and handling events. Practically, this may vary across implementations and be fairly different between simulations and real-world executions. Specifically, defining how the control loop is scheduled may lead to different emerging behaviours of the overall MAS due to which interleaving of agent’s control loops is allowed by the implementation. On the one side, in simulations, the simulator is in charge of (reproducibly) virtualizing the flow of time. In this case, ‘executing the BDI system’ means to advance the simulation time while deciding which agent(s) shall execute their control-loop phases next. The goal here is emulate concurrency, while keeping the execution of the agents deterministic and reproducible. On the other side, in real-world deployments, the agents are executed by the operating system, the flow of time corresponds to the real-world time, and the agents are scheduled by the operating system’s scheduler, which may introduce non-determinism in the execution of the agents. In this case, the goal is to balance agents execution onto the available computational resources, and to exploit parallelism accordingly.
@@ 3 BDI Agents Simulation Methods
This seamless switching is the primary goal for our work [DBLP:conf/dsrt/Baiardi24]. In fact, keeping the same agent specification ensures that simulation results are not artefacts induced by the software translation into a simulation-compatible model. Additionally, if the specification is the same, issues discovered at deployment time that were not evident during testing will be at most imputable to the abstraction gap of the environment model. A recent work has tackled this problem for non-BDI agents [Briola2025], by developing a simulation environment for the JADE MAS platform [DBLP:books/wi/BellifemineCG07]. The work focuses on achieving portability of the same agent behaviour between simulation and real-world execution, by implementing a simulation environment that mimics the JADE run-time environment. We exclude this work from our comparison, as JADE agents are not BDI-based.
@@ 4.1 Portable Agent Specifications
Platform-specific implementations – at least, one for real-world deployments and another for simulation – must be provided for such a platform-agnostic API. We refer to these implementations as the interpreters of the BDI agent specification, which are responsible for executing the agent code by relying on the underlying platform functionalities. The two interpreters may of course differ in their implementations, yet both should guarantee the AgentSpeak(L) semantics is preserved. In this regard, a good practice is to let the two interpreters share as much code as possible: ideally implementing the agent’s control loop in a shared library, which allows for plugging in different platform-specific functionalities.
@@
Agent behaviour may depend on the passage of time. Hence, agents may need to observe the current time, usually through the system’s clock in a real-world deployment. In simulation, however, the flow of time is virtualized and controlled by the simulator, and it usually does not correspond to real-world time. Thus, access to time-dependent functionalities (e.g., timestamps, timeouts, delays) must not be implemented by accessing system-level time primitives directly. Accessing the system’s clock, or using sleep functions must be avoided in the agent code, and instead the environment API must provide the necessary functionalities to access the current time or to schedule timeouts and delays. It will then be the responsibility of the integration with the simulation platform to provide a proper implementation of such functionalities upon the simulator’s notion of time, guaranteeing semantic correspondence with the real-world counterpart.
@@
Agents in a MAS often need to communicate with each other. To support inter-agent communication, the environment API must provide functionalities to send and receive messages. Real-world (distributed) deployments rely on network protocols (e.g., HTTP, MQTT, XMPP, etc.), and although in principle they can be used in simulation as well, a typical choice driven by performance and improved experimental control is to abstract them away and provide direct control over relevant control variables (latency, message/package loss, retransmissions). Typically, in MAS the environment acts as a relay for messages, and messages might be subject to some sort of propagation mechanism. The way messages are propagated may depend on the scenario, for instance, taking into account the environment topology by limiting the delivery of messages to reachable destinations, or introducing some form of delay or loss in the delivery of messages to better match with the expected real-world environment dynamics and stress the properties of the system under different conditions. Again, failing to abstract communication functionalities would make impossible to simulate the system, as the simulator would need to rely on the underlying network stack to deliver messages between agents, which would make the simulation results unreliable and non-reproducible.
@@ 4.4 Tool Requirements for the Integration
Following the key features outlined in the sections above, the tools to be adopted for the integration must provide:
@@
  • simulation model featuring locality, position, and communication channels, to support simulating agents situated in physical or logical spaces;
@@ 5.2 BDI MAS in Alchemist
Each agent’s lifecycle phases (i.e., sense, deliberate, and act) are mapped onto separate DES events (Alchemist’s Reactions). As introduced in Section˜5.1, each such event has a duration modelled through a TimeDistribution. To guarantee that each agent’s control loop phases are executed in order, we implemented a custom TimeDistribution (JaktaTimeDistribution) composed of three sub-TimeDistributions, one for each phase of the control loop. JaktaTimeDistribution maintains an internal state which tracks the current phase of the agent, and executes the next phase according to its configured TimeDistribution. This design allows us to flexibly support different granularities:
@@
Agents can reason on the current device status, as information stored inside Nodes’ is reified as perceived knowledge from JaKtA agents. Information inside nodes is stored as key-value pairs (Molecules) and associated Concentrations (c.f. Section˜5.1), that agents can manipulate through actions. Changes are then treated by the simulator as “regular” environment modifications, consequent to simulated events.
@@ 6.3 Experimental Setup
We ensured code portability by enforcing that all interactions with the environment are done through the JaKtA environment API, which we implemented both for the real-world deployment and for the Alchemist simulator.
6.3.1 Simulated System Deployment on the Alchemist Simulator
We use the simulator to compare the execution of the same MAS logic with different granularities, namely, we consider AMA, ACLI, and ACLP (cf. Section˜4.3). We expect that coarser-grained granularities show better system performance than more finer-grained (in our case, better formation maintenance). The experiment is designed to show that such effect is a consequence of the over-simplification of the BDI program and of its execution using coarse-grained simulation.
@@ 6.3.1 Simulated System Deployment on the Alchemist Simulator
6.3.2 Real-World System Deployment as a Concurrent System
@@ 6.4 Results
Results confirm the expectations that when simulating coarse-grained granularity levels (namely, AMA and ACLI) the agents are able to maintain the formation, while the fine-grained ACLP granularity level reveals that this result is an artefact of the over-simplification of the BDI programs. On a real UAV-programming workflow, this would have revealed a flaw in the agent program, or highlighted the need for higher responsiveness (reasoning cycle frequency) to react to the dynamics of the environment.
6.4.1 Analysis of the Alchemist Simulated System Deployment
@@
6.4.2 Analysis of the Real-World Concurrent System Deployment
@@ 8.1 Achievement of the work objectives
A BDI framework must support modular execution models, abstracting platform-specific functionalities and encapsulating them into a deployment-agnostic API, that allows external scheduling of agents’ reasoning cycles. The framework should maintain a clean separation between agent logic and execution infrastructure, allowing the same specification to run both in simulation and in the real-world target environment without modifications. Implementations should interpret the same agent logic differently. Core to the simulated deployment is the ability to execute BDI agents deterministically, to ensure reproducibility. The same is not true in real-world deployments where concurrency and asynchrony are inherent, and sometimes even desirable and exploitable. Therefore, the deployment-agnostic API should virtualise functionalities covering all potential sources of non-determinism, such as random number generation and external events.
@@
Our integration of JaKtA and Alchemist demonstrates that a BDI framework can be embedded within a DES engine by treating agent execution as a series of simulation-driven events. By leveraging modularity in both technologies, we enable BDI agents to operate in a simulated environment without requiring code modifications. The environment is modeled within the simulator, preserving interactions with agents while ensuring portability. The proposed approach ensures that the same BDI specification can be tested in simulation and seamlessly deployed in the real-world environment (e.g., on distributed computing devices) reducing development overhead and minimizing the reality gap.
@@ Acknowledgements
This work has been partially supported by: “WOOD4.0 - Woodworking Machines for Industry 4.0”, (CUP E69J22007520009) Emilia-Romagna regional project, call 2022, art. 6 L.R. N. 14/2014; and by the projects funded by the European Commission under the NextGenerationEU programme: PNRR – M4C2 – Investment 1.3, Partenariato Esteso PE00000013 – “FAIR—Future Artificial Intelligence Research” – Spoke 8 “Pervasive AI” and PNRR M4C2, Investment 3.3 (DM 352/2022) in collaboration with AUSL Romagna - CUP J33C22001400009 “Digital Twins Ecosystems for the clinical, strategic and process governance in Healthcare”.